OpenStudy (anonymous):

The number of pets owned by eight classmates is shown below: 1, 3, 2, 2, 1, 3, 3, 1 Part A: What is the mean of the data? Show your work. (4 points) Part B: Use your answer from Part A to calculate the mean absolute deviation for the data. Show your work. (6 points)

OpenStudy (anonymous):

I already finished part A

OpenStudy (anonymous):

@Michele_Laino @geny55 @welshfella @LegendarySadist

OpenStudy (anonymous):

i got 2 for the absolute deviation is that correct please help?

OpenStudy (anonymous):

@skullpatrol @ali2x2 @IrishBoy123

OpenStudy (anonymous):

@Michele_Laino

OpenStudy (anonymous):

@Astrophysics

OpenStudy (michele_laino):

first step: the requested mean "m" is given by the subsequent computation: \[\Large m = \frac{{1 + 3 + 2 + 2 + 1 + 3 + 3 + 1}}{8} = ...\]

OpenStudy (anonymous):

16/8= 2

OpenStudy (michele_laino):

correct!

OpenStudy (michele_laino):

second step:

OpenStudy (michele_laino):

for each value v, you have to compute the difference: v-2 and take the result without the sign

OpenStudy (michele_laino):

namely, for example: 1-2= -1, so you take 1 3-2= 1, so you take 1 please continue

OpenStudy (anonymous):

1-2+3-2+2-2+2-2+1-2+3-2+3-2+1-2 so 1,1,0,0,1,1,1,1

OpenStudy (anonymous):

6

OpenStudy (anonymous):

6/8

OpenStudy (michele_laino):

the sum of the absolute deviations is 1+1+1+1+1+1=6 correct and the requested mean is: 6/8=...?

OpenStudy (michele_laino):

since we have 8 entries

OpenStudy (anonymous):

erm 1.2?

OpenStudy (michele_laino):

6/8=3/4=...?

OpenStudy (anonymous):

0.75

OpenStudy (michele_laino):

correct!

OpenStudy (anonymous):

Thank you for making it so much easier to understand (: @Michele_Laino I love your help

OpenStudy (michele_laino):

:)